2x+3y=12;5x-3y=9(by substituion method)

Answer by kadash12(6)

2x+3y=12;5x-3y=9
Start by solving one function equal to y. We'll use 2x+3y=12 this time.
2x+3y=12
-2x -2x
3y=-2x+12
---------= y=-2/3x+4
3
Now since y=-2/3x+4 then 5x-3y=9
^--this y would also equal to -2/3x+4 so sub in.
5x-3(-2/3x+4)=9 Distribute
5x+2x-12=9 Collect like terms
7x=21 Divide
x=3 Now to get y you sub in 3 to x.
2(3)+3y=12
6+3y=12 collect like terms.
3y=6 divide
y=2